OpenStudy (anonymous):

10^(2x+4)=3 what is x

OpenStudy (anonymous):

You can take the log in base ten of both sides

OpenStudy (anonymous):

1.7614

OpenStudy (anonymous):

Take the common logarithm (base 10) of both sides:\[\log10^{2x+4}=\log3 \implies (2x+4)\log(10)=\log3 \implies 2x+4 = \log3 \implies x = ..?\] Can you solve for x from here? @Breno